OpenStudy (anonymous):

An isotope of an element with a relative atomic mass of 19 amu and 10 neutrons has a charge of 1-. How many electrons does it have?

sam (.sam.):

amu= proton + neutron 19= proton + 10 proton=9 -------------------------------- We know that proton = electron Electron =9 But it has an extra charge of -1, so, Electron =10

OpenStudy (unklerhaukus):

\[^A_Z\text E^{q}_N\] \[A=Z+N\]\[Z=e+q\] \[^{19}_Z\text E^{-1}_{10}\] \[19=Z+10\qquad\rightarrow Z=19-10\]\[Z=e-1\qquad\rightarrow e=Z+1\]